The perimeter of a college basketball court is 108 meters abd the lenght is twice as long as the width. What are the lenght and width>
Answer by RAY100(1637)

per = 2L+2w
but L=2w
and per = 108
subst
108= 2(2w) + 2w = 6w
divide both sides by 6

18=w

L=2w =2*18 = 36
check
108=2(18)+2(36)= 108 ok